OpenStudy (anonymous):

Limit X-->1 f(x) = 2x^2-x-1/3x^2+x-4?

OpenStudy (anonymous):

when i worked it out i got 3/7 ???

OpenStudy (anonymous):

factor and cancel, replace \(x\) by \(1\)

OpenStudy (anonymous):

\lim_{x \rightarrow 1}\frac{2x^2-x-1)}{3x^2+x-4}\] Using l'Hopital I also got 3/7

OpenStudy (anonymous):

cool thanks

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\frac{2x^2-x-1}{3x^2+x-4}\]

OpenStudy (anonymous):

\[\frac{(x-1)(2x+1)}{(x-1)(3x+4)}=\frac{2x+1}{3x+4}\]

OpenStudy (anonymous):

1 gets you \(\frac{3}{7}\)

OpenStudy (anonymous):

yes factor top and bottom then cancel the x-1. Replace x by in 1 in 2x+1/3x+4 should equal 3/7